Find the slope of the line passing through the points (2,-4) and (7,5)
Answer by lwsshak3(11628)

Find the slope of the line passing through the points (2,-4) and (7,5)
slope of line=∆y/∆x=change in y/change in x=rise/run=(5-(-4))/(7-2)=9/5